Can an umbrella 28 cm long be packed in the bottom of a suitcase 24 cm long by 16 cm?
Answer by Alan3354(69443)

Can an umbrella 28 cm long be packed in the bottom of a suitcase 24 cm long by 16 cm?
-----------
The diagonal = sqrt(24^2 + 16^2)
= sqrt(576 + 256)
= sqrt(832)
= 8*sqrt(13)
=~ 28.844 cm, so it will fit.
The umbrella is only 11 inches long, so what good will it do? Is it for someone's ferret?